OpenStudy (aonz):

Find the gradients of the lines that pass through the point (1, 7) and are tangent to the parabola y = (2 − x)(1 + 3x)

OpenStudy (aonz):

@oldrin.bataku help please

OpenStudy (aonz):

expanded -3x^2 + 5x + 2

OpenStudy (amistre64):

define the derivative now

OpenStudy (aonz):

-6x + 5

OpenStudy (amistre64):

then that will define the slope of the tangent line for some x=a; attach that to the given point

OpenStudy (aonz):

m = -1?

OpenStudy (amistre64):

no, m=(-6a+5) attached to the point 1,7

OpenStudy (aonz):

? what do you mean attach it?

OpenStudy (amistre64):

do you recall how to form a line given a slope and a point?

OpenStudy (aonz):

y-y1 = m(x-x1)

OpenStudy (amistre64):

correct, use that ....

OpenStudy (aonz):

y-7 = (5-6a)(x-1)

OpenStudy (amistre64):

now, for some point (a,b); the parabola and the line equal ... so equate them.

OpenStudy (aonz):

x^2 - 10x + 6ax - 6a - 5 = 0

OpenStudy (amistre64):

this is where my thought goes; for some x=a -3a^2 + 5a + 2 = (-6a+5)a + 6a +2

OpenStudy (amistre64):

solve for a

OpenStudy (aonz):

using B^2 - 4ac right?

OpenStudy (amistre64):

if you want to use the quadratic formula, then thats fine.

OpenStudy (aonz):

i have gotten 3a^2 - 8a + 10 I think i went wrong somewhere

OpenStudy (amistre64):

-3a^2 + 5a + 2 = (-6a+5)a + 6a +2 -3a^2 + 5a = (-6a+5)a + 6a -3a^2 = -6a^2+5a + a 3a^2 -6a = 0 3a(a -2) = 0, when a=0, or a=2

OpenStudy (aonz):

ok now we sub it into y=mx+b? or something like that?

OpenStudy (amistre64):

well, we already have the line equation; we just need to fill in our a values y = (5-6a)(x-1)+7 y1 = 5(x-1)+7 y2 = -7(x-1)+7 http://www.wolframalpha.com/input/?i=y%3D-3x%5E2+%2B+5x+%2B+2%2C+y-7+%3D+%285%29%28x-1%29%2C+y-7+%3D+%285-6%282%29%29%28x-1%29

OpenStudy (aonz):

So gradients would be 5 and -7?

OpenStudy (amistre64):

yep

OpenStudy (aonz):

thanks for your help

OpenStudy (amistre64):

youre welcome